OpenStudy (anonymous):

A kayaker can paddle 12 mi in 2 h moving with the river current. Paddling at the same pace, the trip back against takes 4 h. Assume that the river current is constant. Find what the kayakers speed would be in still water.

OpenStudy (anonymous):

Easy. If the kayaker paddles 12 mi in 2h he will paddle 24 in 4 hours.

OpenStudy (redheadangel):

Yes, but it asked for in STILL water (other than that your reasoning is correct)

OpenStudy (anonymous):

I see...

OpenStudy (amistre64):

since the distance is the same each time, equate distance with speed and time

OpenStudy (anonymous):

ok do you think you could set it up for me please

OpenStudy (amistre64):

hmm, give me your attempt first. how are distance speed and time related?

OpenStudy (anonymous):

d= R*T

OpenStudy (amistre64):

good, lets use distance = speed * time since the distance is the same we can say: distance = (river + boat) * time1 distance = (river - boat) * time2 something to that effect

OpenStudy (amistre64):

we have a distance given to us in miles for a time of 2 hours 12miles = (river + boat) * 2 hours if we assume that the distance is 12 we can compare it with 12miles = (river - boat) * 4 hours

OpenStudy (anonymous):

ok I get so like r+b=6 r-b=3

OpenStudy (amistre64):

good, its a system of eqautions ...

OpenStudy (amistre64):

if we add them as is we eliminate the boat :) 2r = 9 when r=9/2 so we know the river spped

OpenStudy (anonymous):

did I set my equation up right

OpenStudy (amistre64):

yeah, you did great

OpenStudy (anonymous):

ok give me a sec and tell me if my answers right

OpenStudy (anonymous):

ok for r I got 4.5 and b I got 1.5

OpenStudy (amistre64):

my only concern with this thought is that we are not really given an actual distance, just a rate of speed. if the distance was 12, then the river is 4.5 and the boat is 1.5 i agree.

OpenStudy (anonymous):

thankyou very much

OpenStudy (amistre64):

youre welcome